# Code-Driven Law NO, Normware SI!

### v1

@@ 3.1 Directives concerning regulation
In the most intuitive interpretation, “normware” refers to computational artifacts specifying norms, ie. intelligible directives concerning regulation of behaviour. Common computational artifacts going under this definition would be access control and usage control policies, machine-readable data-sharing agreements, digital consents, policies for digital infrastructure regulating eg. routing, or data-sharing across domains, as well as behavioural policies (scripts) for computational agents. Complementary to these more operational contributions, there is a large body of literature on normative systems, consisting of a vast number of contributions about deontic logic, i.e. the logic of obligation, prohibition, and permission (see the overviews given in [18, 19]), and fewer on other conceptual frameworks (eg. Hohfeld’s primitive normative concepts [20, 21, 22, 23, 24, 25]). Yet, it is crucial to observe that any computer program has inherently a “normative” interpretation, even without referring to explicit normative concepts, or used explicitly for normative purposes.
@@
The most traditional form of programming artifacts is that of imperative programs. Imperative programs consist of sequences of instructions, ie. actions that ought to be performed by the machine. Even machine code, used to interact with hardware, is generally expressed in imperative terms. However, if the engineering objective for normware artifacts is to achieve some form of legal isomorphism [14] (or, more reasonably, a weaker version of it, ie. targeting not the text of normative source, but a certain interpretation of the text, expressed by some expert), imperative programs cannot specify the diversity of institutional constructs that can be found in normative sources (contracts, legislation, …). For instance, concepts as prohibitions cannot be directly operationalized one-on-one to imperative programs, as we can define only what the machine has to perform, positively. [^6] It is worth to note that Solidity, the most used language today to specify smart contracts, is an imperative programming language. This entails that, from a designer perspective, the main value of contemporary smart contract technology lies in providing by design relevant non-functional requirements (eg. immutability) via distributed ledgers and decentralized execution, rather than a sound representational model to specify contractual mechanisms.
@@
Several authors in the literature (see eg. [26, 27]) have observed that declarative programs may be a more suitable tool for representing norms, as norms may be specified in terms of constraints that ought to be satisfied (sometimes called “goals”). Dedicated solvers can then be used to automate the process of finding solutions (eg. behaviours) that satisfy these constraints. In most cases, goals can also be defined negatively, with the meaning “this should not occur”, where this—simplifying—is some abstract or concrete situation (eg. object or actor being in a certain state, an actor performing a certain action). More complex notations take also into account preferences (also named weak goals, or soft constraints), ie. the fact that goals may have a relative (and conditional) priority. Computational methods have been introduced to automatically derive the most preferred goals, under certain assumptions. This is particularly relevant in domains in which goals can be conflicting, as priorititization enables or at least facilitates conflict resolution.
@@ 3.2 Directives concerning qualification
As soon as we climb the ladder of abstraction we are however confronted with the classic problems of semantics: What do we mean by this? (for language-mediated situations), correlative to What is this? (for concrete, direct situations). Part of these wide issues in philosophy and in symbolic AI goes under the name of qualification problem [28], concerning the adequate expression of the preconditions required for an action (here, a certain naming) to be applicable. Interestingly, law attempts to partially solve similar challenges by means of constitutive rules defining within the normative sources what count-as what, or, complementarily, via procedural heuristics like evidential rules emerging from case law.
Fortunately, being in the computational domain provides us with a privileged (but also illusive) starting point to tackle this problem, in comparison to a human-centred social domain. This is because hardware manufacturers physically hard-code the (local) semantics of primitive operations, consisting of logical and arithmetic operations, and specific instructions controlling memory or some input/output modules. As long as all possible “this” for our domain of concerns can be referred back to these physical primitives, or their logical counterparts, the qualification problem can be solved. However, because in any practical application we need to go beyond the computational boundary (that is, there are always elements referring to or depending on what is outside the computational system), we cannot truly bypass the qualification problem, and this entails that we need terminological specifications provided by humans to scope more adequately what is being regulated. These artifacts are also part of normware and aim to capture correct terminology, rather than correct (or wrong) behaviour.
@@ 3.3 Directives concerning expectations
As a discipline, knowledge engineering targets more widely ontologies: shared conceptualizations of a target domain expressed in a machine-processable form [29]. These conceptualization artifacts contain terminological knowledge, but they also include relationships attributed (in most cases by humans) to the target domain: eg. taxonomies, partonomies, causal dependencies and logical constraints. Independently from the specific representational model associated to the various solutions introduced in the literature and by industry, all these artifacts aim to reify expectations in symbolic forms. This was deemed relevant for computational regulative purposes, as most normative reasoning tasks require some form of world knowledge [30], eg. to elaborate on causal chainings for responsibility attribution, to compare the impact of action vs omission of action, to evaluate foreseeability of events. Besides, this elaboration concerns not only physical mechanisms/outcomes but also symbolic ones, as for instance with institutional powers, identifying and handling actions whose performance carries institutional meaning. Unfortunately, it is also well-known that computational ontologies and more generally purely symbolic methods suffer from serious limitations (e.g. the infamous knowledge-acquisition bottleneck, the difficulty of alignment across ontologies, the lack of symbol grounding [31], the ramification and qualification problems, and so on), therefore ontologies (or similar types of technologies) cannot be deemed to be a sufficient mean.
Looking at a higher abstraction level, ontological artifacts are meant to capture what is to be expected, or is at least possible in the target domain. Indeed, beside normativity, the “norm” in normware refers also to a normality dimension. Statistical methods are particularly suited to identify what is “normal” in a certain sample space, and the same applies with statistical machine-learning: latent spaces captured in machine-learning models are non-intelligible expression of regularities induced from observations. For instance, a “cat” classifier embeds some model of a prototypical cat, [^7] To be more precise, machine-learning models are traditionally divided in discriminative or generative, depending on the information principle applied for selecting the output. Only generative models are seen as maintaining a prototype concept of cat, as they allow us to produce a “cat” output if required so. Yet, discriminative models still maintain a prototype of an evidential process meant to reach a conclusion on whether a given input is a cat. time-series predictors embeds models of temporal patterns, word-embeddings reify expectations of proximity between “concepts”. May these artifacts have also a role within a computational framework centred on normware? To answer to this question, we need to take a different route: we need to abstract the fact that there is a designer that intentionally specified the device; in other words, we need to look at artifacts in functional terms.
@@ 3.4 Devices intended to regulate
All designed artifacts can be interpreted just by knowing the function for which they have been designed. A door regulates the entrance to a certain building or room. A semaphore signals to vehicles, bikes and pedestrians whether they are allowed or prohibited to cross. Suppose the door to be also electronically controlled. In both cases (semaphore, and controlled door), we do not have access to the inner decision-making mechanism that brought to the current output state, but, as far as we can see this output (eg. the door being closed, the light being red), and we know to what is referring to (eg. allowing a certain action or not), we are able to interpret it, adapting our behaviour accordingly. The case of the door goes even further, as the device intervenes in the physical environment disabling our action without requiring our understanding. These examples support the idea that black-boxes as machine-learning models are artifacts expressing some form of normality/normativity, and as such they can be thought as normware artifacts too. In other words, under the lens of normware, there is no functional difference between code-driven and data-driven law. [^8] The distinction is in terms of implementation. In code-driven law, the underlying directives supporting a certain conclusion are explicit and made intelligible by relying on symbols whose meaning was (supposedly) stipulated at a certain moment, and expressed in the form of program; in data-driven law, these directives are tacit, expressed in the form of eg. parameters of a neural network. Both forms of computation are meant to regulate behaviours, qualifications, and expectations. For instance, a model trained via reinforcement learning will provide as output actions that have to be performed conditionally to the observed input, a classifier will provide labels to input objects, a temporal predictor projections in the future of given states, an autoencoder a latent model of the domain from which we can identify abnormal instances, and so on.
@@ 4.1 Regulation as control
Institutional systems since ancient times are based on some assumption of intentionality (eg. the cases in [32, p. 102]). The behaviour of intentional agents, by definition, can be interpreted in terms of purpose, motives, and motivations. In many cases, however, ‘weaker’ interpretations of action may also apply, which do not require the existence of a mental state associated with the intent, but still maintain a purposive character. For instance, the notion of control system has a primary role in engineering disciplines, and issues of regulation (eg. homeostasis) are central to the study of biological systems. A crucial difference between control systems and agents is that the latter are deemed to be able to generate their own goals autonomously (albeit with various levels of autonomy), while for the former, goals have been hard-coded by design (top-down), or by selection through some adaptive, possibly evolutionary process (bottom-up). The goal of an entity is then, in an ecological perspective, the function of the entity within the environmental niche in which it is embedded, and transcends the entity itself. This is clear for designed artifacts, but natural entities also fall under similar considerations without taking a creationist stance (cf. the concept of teleonomy introduced by Monod [33] as counterpart to teleology). Whether artificial or natural, designed or emergent, what counts in control is the existence of some reference (the target of control) indicating a “point” or “region” in the space of possible behaviours or states of the system, which the entity is set to either achieve or avoid (the direction of control). [^9] More complex systems may additionally exhibit the direction “to avoid to control” (ie. maintaining no-direction) with respect to the target, cf. the concept of strong permission in the normative systems literature [34, 35]. By defining directives by means of this control signature (target and direction), any regulative mechanisms can be abstracted from its implementation.
@@ 4.2 Indeterminacy of references and directives
The legal activity and jurisprudential debates make clear that, for socially relevant phenomena, references cannot be defined once and for all, nor expressed entirely with a single artifact, and not even taking into account all normative artifacts (pace strong legal positivism). A sounder stance is to see these situational “points” (eg. a fairness point) as mere virtual placeholders for setting up socio-technical arrangements, contextually and historically dependent [36]. The correct functioning of these arrangements is enabled only by some common understanding or behavioural coordination amongst participants, which they have agreed, or accepted, or at least not contrasted. This acknowledgement entails that we cannot rely on the assumption that there is only one correct way to model (symbolically or sub-symbolically) what a “cat” is, in all circumstances; that all what a contract means can be reified into a single artifact; that there exists an abstract fairness point, that can be discovered once and for all by some mathematical formula.
@@
Indeterminacy of references entails indeterminacy of directives, but indeterminacy of directives may also occur because of antinomies, ie. when two or more conflicting directives exist. Legal systems generally introduce formal criteria to identify the relative priority of norms (eg. lex specialis, superior, and posterior), but it is generally in the hard-cases that one can see that formal criteria are accompanied by underlying/implicit preferential structures, eventually embodied by the judges (“voices of the law”), and sometimes expressed as values (see eg. [37] in the case-based reasoning literature), that are meant to balance or complement formal mechanisms (or their absence) to supposedly obtain a more just result. At a fundamental level, the primary source of these indeterminacies are the different positions, perceptions, knowledge, expectations, motivations, and preferences of the various participants to the social gathering, and this diversity in perspectives brings naturally to different qualifications. [^10] See for instance the argumentative theory of reasoning [38], humans capture from the world what it is deemed relevant to advance their interests, they do not strive for truth in absolute sense. Distinct selections lead consequently to distinct qualifications.
@@
All these observations apply similarly to components of distributed computational systems. Computational modules with the same function may be given different inputs (eg. two similar sensors placed in different positions), and/or defined by distinct processes (either designed or trained) striving for different goals or value structures, and for this reason may produce conflicting output. Moreover, humans (with all their heterogeneity) intervene in the computational loop, for instance interacting via some interface with the computational system, or possibly feeding it with additional data and/or modifying its code. Conflicting directives cannot but be deemed systematic.
@@ 4.3 Coordination and conflict resolution
Taken as an atomic component, a piece of normware functions as a coordination mechanism, just like a semaphore. The concern of such coordination may be prescriptive, with normware artifacts meant to regulate other modules in a certain way (w.r.t. behaviour, or terminology); or descriptive, in the sense that a certain coordination function is observed or ascribed to a certain component (in the form of qualifications, or expectations). However, the introduction of the norwmare level of computation is purposely meant to go beyond the single individual artifact, and to acknowledge the “ecological” dimension of computation, ie. the presence and entrenchment of several concurrent—possibly conflicting—coordination mechanisms within the same system. The ecological dimension becomes also evident when there are frictions between descriptive and prescriptive perspectives upon the same entity.
@@
Failures of the second type are purely internal to the system, and require the inhibition or the removal of a number of components to restore directivity. These conflicts concern the viability of the system. A system which is unable to restore directivity carries the risks of becoming idle, which often does not coincide with begin neutral between the two directions, but de facto taking one of the two by default. For instance, in a context in which both duty of $A$ and prohibition of $A$ are issued, not doing anything means fulfilling the second.
Note also that it is not necessary for failures to occur to trigger a system’s response. Expectations can have a crucial role for anticipating conflicts, promoting preventive or preparatory behaviour. The various risk strategies can for instance be thought as meta-directives dedicated to anticipatory functions, capturing the entrenchment between expectations and references targeted by regulations.
@@
The maintenance of processes of determination of references, and of resolution of conflicts requires in itself some reference/directive. This requirement links to the procedural dimension of law. Conflict resolution directives can in principle also be contested, and dynamically changed. In the human realm, any contract needs to comply with contract law, which defines some procedural, formal requirements that make a contract valid. Contract law in itself has been issued by some legislative body, whose members—supposing a democratic system—were elected, with voting mechanisms defined in turn by some previous legislative body. This example is relevant here for two reasons. First, it shows that determination/resolution of conflicts can be proceduralized in modular forms within the same normative system depending on the current institutional architecture: eg. aggregation of preferences by majority, formal criteria, intervention of humans at certain decision points (eg. contract parties, judges, legislators, voters). Second, it raises the question of how a normative system should originate in order to unfold.

### v2

@@ 3.1 Directives concerning regulation
In the most intuitive interpretation, “normware” refers to computational artifacts specifying norms, i.e. intelligible directives concerning regulation of behaviour. Common computational artifacts going under this definition would be for instance access control and usage control policies, machine-readable data-sharing agreements, digital consents, policies for digital infrastructure regulating e.g. routing, or data-sharing across domains, as well as behavioural policies for computational agents. Complementary to these more operational contributions, there is a large body of literature on normative systems, consisting of a vast number of contributions on deontic logic, i.e. the logic of obligation, prohibition, and permission (see the overviews given in [18, 19]), and fewer on alternative conceptual frameworks (e.g. Hohfeld’s primitive normative concepts [20, 21, 22, 23, 24, 25]). Yet, even without referring to explicit normative concepts, or used explicitly for normative purposes, any computer program has inherently a “normative” interpretation.
@@
The most traditional form of programming artifacts is that of imperative programs. Imperative programs consist of sequences of instructions, i.e. actions that ought to be performed by the machine. Even machine code, used to interact with hardware, is generally expressed in imperative terms. However, if the engineering objective for normware artifacts is to achieve some form of legal isomorphism [14] (or, more reasonably, a weaker version of it, i.e. targeting not the text of normative source, but a certain interpretation of the text, expressed by some expert), imperative programs cannot specify the diversity of institutional constructs that can be found in normative sources. For instance, concepts as prohibitions cannot be directly operationalized one-on-one to imperative programs, as we can define only what the machine has to perform, positively. [^6] It is worth to note that Solidity, the most used language today to specify smart contracts, is an imperative programming language. This entails that, from a designer perspective, the main value of contemporary smart contract technology lies in providing by design relevant non-functional requirements (e.g. immutability) via distributed ledgers and decentralized execution, rather than a sound representational model to specify contractual mechanisms.
@@
Several authors in the literature (see e.g. [26, 27]) have observed that declarative programs may be a more suitable tool for representing norms, as norms may be specified in terms of constraints that ought to be satisfied (sometimes called “goals”). Dedicated solvers can then be used to automate the process of finding solutions (e.g. behaviours) that satisfy these constraints. In most cases, goals can also be defined negatively, with the meaning “this should not occur”, where this—simplifying—is some abstract or concrete situation (e.g. object or actor being in a certain state, an actor performing a certain action). More complex notations take also into account preferences (also named weak goals, or soft constraints), i.e. the fact that goals may have a relative (and conditional) priority. Several computational methods have then been introduced to automatically derive the most preferred goals, under certain assumptions. A declarative approach is particularly relevant in domains in which goals can be conflicting, as priorititization enables or at least facilitates conflict resolution.
@@ 3.2 Directives concerning qualification
As soon as we climb the ladder of abstraction we are however confronted with the classic problems of semantics: What do we mean by this? (for language-mediated situations), correlative to What is this? (for concrete, direct situations). Part of these wide issues in philosophy and in symbolic AI goes under the name of qualification problem [28], concerning more generally the adequate expression of the preconditions required for an action (here, a certain naming) to be applicable. Interestingly, law attempts to partially solve similar challenges by means of constitutive rules defining within the normative sources what count-as what, or, complementarily, via procedural heuristics like evidential rules emerging from case law.
Fortunately, being in the computational domain provides us with a privileged (but also illusive) starting point to tackle this problem, in comparison to a human-centred social domain. This is because hardware manufacturers physically hard-code the (local) semantics of primitive operations, consisting of logical and arithmetic operations, and specific instructions controlling memory or some input/output modules. As long as all possible “this” for our domain of concerns can be referred back to these physical primitives, or their logical counterparts, the qualification problem can be solved. However, because in any practical application we need to go beyond the computational boundary (there are always elements referring to or depending on what is outside the computational system), we cannot truly bypass the qualification problem, and this entails that we need terminological specifications provided by humans to scope more adequately what is being regulated. These artifacts are also part of normware and aim to capture correct terminology, rather than correct (or wrong) behaviour.
@@ 3.3 Directives concerning expectations
Terminology is however not sufficient in itself. As a discipline, knowledge engineering emerged as targeting more widely ontologies: shared conceptualizations of a target domain expressed in a machine-processable form [29]. These conceptualization artifacts contain terminological knowledge, but they also include relationships attributed (in most cases by humans) to the target domain: e.g. taxonomies, partonomies, causal dependencies and logical constraints. Independently from the specific representational model associated to the various solutions introduced in the literature and by industry, all these artifacts aim to reify expectations in symbolic forms. This was deemed relevant for computational regulative purposes, as most normative reasoning tasks require some form of world knowledge [30], e.g. to elaborate on causal chainings for responsibility attribution, to compare the impact of action vs omission of action, to evaluate foreseeability of events. Besides, this elaboration concerns not only physical mechanisms/outcomes but also symbolic ones, as for instance with institutional powers, specifying actions whose performance carries institutional meaning. Unfortunately, it is also well-known that computational ontologies and more generally purely symbolic methods suffer from serious limitations (e.g. the infamous knowledge-acquisition bottleneck, the difficulty of alignment across ontologies, the lack of symbol grounding [31], the ramification and qualification problems, and so on), therefore ontologies (or similar types of technologies) cannot be deemed to be a sufficient mean.
Now, ontological artifacts are meant to capture what is to be expected, or is at least possible in the target domain. Indeed, beside normativity, the “norm” in normware captures also a normality dimension. Interestingly, statistical methods are particularly suited to identify what is “normal” in a certain sample space, and the same applies with statistical machine-learning: latent spaces captured in machine-learning models are non-intelligible expression of regularities induced from observations. For instance, a “cat” classifier embeds some model of a prototypical cat, [^7] To be more precise, machine-learning models are traditionally divided in discriminative or generative, depending on the information principle applied for selecting the output. Only generative models are seen as maintaining a prototype concept of cat, as they allow us to produce a “cat” output if required so. Yet, discriminative models can still be seen as reifying an evidential process meant to reach a conclusion on whether a given input is a cat. time-series predictors embed models of temporal patterns, word-embeddings reify expectations of proximity between “concepts”, large-language models (LLMs) predict next words given an initial textual context. May these artifacts have also a role within a computational framework centred on normware?
@@ 3.4 Devices intended to regulate
In general, designed artifacts can be interpreted by knowing the function for which they have been designed. A door regulates the entrance to a certain building or room. A semaphore signals to vehicles, bikes or pedestrians whether they are allowed or prohibited to cross. Suppose the door to be also electronically controlled. In both cases (semaphore, and controlled door), we do not have access to the inner decision-making mechanism that brought to the current output state, but, as far as we can see this output (e.g. the door being closed, the light being red), and we know to what is referring to (e.g. allowing a certain action or not), we are able to interpret it, adapting our behaviour accordingly. The case of the door goes even further, as the device intervenes in the physical environment disabling our action without requiring our understanding. These examples support the idea that black-boxes as machine-learning models are artifacts expressing some form of normality/normativity, and as such they can be thought as normware artifacts too. In other words, under the lens of normware, there is no difference between code-driven and data-driven law. [^8] The distinction is in terms of implementation. In code-driven law, the underlying directives supporting a certain conclusion are explicit and made intelligible by relying on symbols whose meaning was (supposedly) stipulated at a certain moment, and expressed in the form of program; in data-driven law, these directives are tacit, expressed in the form of e.g. parameters of a neural network. Both forms of computation are meant to regulate behaviours, qualifications, and expectations. For instance, a model trained via reinforcement learning will provide as output actions that have to be performed conditionally to the observed input, a classifier will provide labels to input objects, a temporal predictor projections in the future of given states, an autoencoder a latent model of the domain from which e.g. to identify abnormal instances, a large language model a text completing a given prompt, and so on.
@@ 4.1 Regulation as control
Institutional systems since ancient times are based on some assumption of intentionality (e.g. the cases in [32, p. 102]). The behaviour of intentional agents, by definition, can be interpreted in terms of purpose, motives, and motivations. In many cases, however, weaker interpretations of action may also apply, which do not require the existence of a mental state associated with the intent, but still maintain a purposive character. For instance, the notion of control system has a primary role in engineering disciplines, and issues of regulation (e.g. homeostasis) are central to the study of biological systems. A crucial difference between control systems and agents is that the latter are deemed to be able to generate their own goals autonomously (albeit with various levels of autonomy), while for the former, goals have been hard-coded by design (top-down), or by selection through some adaptive, possibly evolutionary process (bottom-up). The goal of an entity is then, in an ecological perspective, the function of the entity within the environmental niche in which it is embedded, and transcends the entity itself. This is clear for designed artifacts, but natural entities also fall under similar considerations without taking a creationist stance (cf. the concept of teleonomy introduced by Monod [33] as counterpart to teleology). Whether artificial or natural, designed or emergent, what counts in control is the existence of some reference (the target of control) indicating a “point” or “region” in the space of possible behaviours or states of the system, which the entity is set to either achieve or avoid (the direction of control). [^9] More complex systems may additionally exhibit the direction “to avoid to control” (i.e. maintaining no-direction) with respect to the target, cf. the concept of strong permission in the normative systems literature [34, 35]. By defining directives by means of this control signature (target and direction), any regulative mechanisms can be abstracted from its implementation.
@@ 4.2 Indeterminacy of references and directives
The legal activity and jurisprudential debates make clear that, for socially relevant phenomena, references cannot be defined once and for all, nor expressed entirely with a single artifact, and not even taking into account all normative artifacts (unless we take a strong legal positivist position). A sounder stance is to see these situational “points” as mere virtual placeholders for setting up and discussing socio-technical arrangements, contextually and historically dependent [36]. The correct functioning of these arrangements is enabled only by some common understanding or behavioural coordination amongst participants, which they have agreed, or accepted, or at least not contrasted. This acknowledgement entails that we cannot rely on the assumption that there is only one correct way to model (symbolically or sub-symbolically) what a “cat” is, in all circumstances; that all what a contract means can be reified into a single artifact; that there exists an abstract fairness point, that can be discovered once and for all by some mathematical formula, and so on.
@@
Indeterminacy of references entails indeterminacy of directives, but indeterminacy of directives may also occur because of antinomies, i.e. when two or more conflicting directives exist. Legal systems generally introduce formal criteria to identify the relative priority of norms (e.g. lex specialis, superior, and posterior), but it is generally in the hard-cases that one can see that formal criteria are accompanied by underlying/implicit preferential structures, eventually embodied by the judges (“voices of the law”), and sometimes expressed as values (see e.g. [37] in the case-based reasoning literature), that are meant to balance or complement formal mechanisms (or their absence) to supposedly obtain a more just result. At a fundamental level, the primary source of these indeterminacies are the different positions, perceptions, knowledge, expectations, motivations, and preferences of the various participants to the social gathering, and this diversity in perspectives brings naturally to different qualifications. [^10] See for instance the argumentative theory of reasoning [38], humans capture from the world what it is deemed relevant to advance their interests, they do not strive for truth in absolute sense. Distinct selections lead consequently to distinct qualifications.
@@
All these observations apply similarly to components of distributed computational systems. Computational modules with the same function may be given different inputs (e.g. two similar sensors placed in different positions), and/or defined by distinct processes (either designed or trained) striving for different goals or value structures, and for this reason may produce conflicting output. Moreover, humans (with all their heterogeneity) intervene in the computational loop, for instance interacting via some interface with the computational system, or possibly feeding it with additional data and/or modifying its code. Conflicting directives cannot but be deemed systematic.
@@ 4.3 Coordination and conflict resolution
Taken as an atomic component, a piece of normware functions as a coordination mechanism, just like a semaphore. The concern of such coordination may be prescriptive, with normware artifacts meant to regulate other modules in a certain way (w.r.t. behaviour, or terminology); or descriptive, in the sense that a certain coordination function is observed or ascribed to a certain component (in the form of qualifications, or expectations). However, the introduction of the norwmare level of computation is purposely meant to go beyond the single individual artifact, and to acknowledge the “ecological” dimension of computation, i.e. the presence and entrenchment of several concurrent—possibly conflicting—coordination mechanisms within the same system. The ecological dimension becomes also evident when there are frictions between descriptive and prescriptive perspectives upon the same entity.
@@
Failures of the second type are purely internal to the system, and require the inhibition or the removal of a number of components to restore directivity. These conflicts concern the viability of the system. A system which is unable to restore directivity carries the risks of becoming idle, which often does not coincide with being neutral towards the two directions, but de facto taking one of the two by default. For instance, in a context in which both a duty of $A$ and a prohibition of $A$ are issued, not doing anything means fulfilling the second directive.
Note also that it is not necessary for failures to occur to trigger a system’s response. Expectations have a crucial role for anticipating conflicts, promoting preventive or preparatory behaviour. The various risk strategies can for instance be thought as meta-directives dedicated to anticipatory functions, capturing the entrenchment between expectations and references targeted by directives.
@@
The maintenance of processes of determination of references, and of resolution of conflicts requires in itself some reference/directive. This requirement links to the procedural dimension of law. In turn, conflict resolution directives can also be contested, and dynamically changed. In the human realm, any contract needs to comply with contract law, which defines some procedural, formal requirements that make a contract valid. Contract law in itself has been issued by some legislative body, whose members—supposing a democratic system—were elected, with voting mechanisms defined in turn by some previous legislative body. The example is relevant for two reasons. First, it shows that determination/resolution of conflicts can be proceduralized in different, modular forms within the same normative system: e.g. aggregation of preferences by majority, formal criteria, intervention of humans at certain decision points (e.g. contract parties, judges, legislators, voters). Second, it raises the question of how a normative system should originate in order to unfold.